The Modern Grocery has cashews that sell for $3.75 a pound and peanuts that sell for $2.75 a pound. How much of each must Albert, the grocer, mix to get 40 pounds of mixture that he can sell for $3.00 per pound? Express the problem as a system of linear equations and solve using the method of your choice.
Answer by checkley77(12569)

3.75x+2.75(40-x)=3.00*40
3.75x+110-2.75x=120
x=120-110
x=10 pounds of cashews is used.
40-10=30 pounds of peanuts are used.
Proof:
3.75*10+2.75*30=120
37.5+82.5=120
120=120